# HydraLoRA: An Asymmetric LoRA Architecture for Efficient Fine-Tuning

### v1

@@ Abstract
Adapting Large Language Models (LLMs) to new tasks through fine-tuning has been made more efficient by the introduction of Parameter-Efficient Fine-Tuning (PEFT) techniques, such as LoRA. However, these methods often underperform compared to full fine-tuning, particularly in scenarios involving complex datasets. This issue becomes even more pronounced in complex domains, highlighting the need for improved PEFT approaches that can achieve better performance. Through a series of experiments, we have uncovered two critical insights that shed light on the training and parameter inefficiency of LoRA. Building on these insights, we have developed HydraLoRA, a LoRA framework with an asymmetric structure that eliminates the need for domain expertise. Our experiments demonstrate that HydraLoRA outperforms other PEFT approaches, even those that rely on domain knowledge during the training and inference phases. Code.
† Equal Contribution. * Corresponding author
@@ 1 Introduction
Large Language Models (LLMs; [lm-4, lm-5, lm-6, llama, llama2, chatgpt, gpt4]) are notably powerful, yet their training involves substantial expense. Adapting a single LLM for multiple downstream applications via fine-tuning has emerged as a prevalent method to cater to specific domain needs, balancing performance with practicality. This approach, however, faces a significant challenge due to the extensive memory and computational resources required for full fine-tuning (FFT), i.e., fine-tuning all billions of parameters. A solution to this has been the development of more selective adaptation techniques, involving modifying only a portion of the parameters or integrating external modules designed for new tasks. Key methodologies in this sphere include LoRA [HuSWALWWC22], Adaptors [RebuffiBV17, HoulsbyGJMLGAG19, MahabadiHR21], and many other variants [LiL20, LesterAC21, DengWHWGSSXH22, HeZMBN22, ZakenGR22], all part of what can be generally termed as Parameter-Efficient Fine-tuning (PEFT). PEFT strategies are characterized by freezing the backbone model parameters while only a minimal number of task-specific parameters are introduced and fine-tuned. This method substantially boosts efficiency in the phases of fine-tuning and subsequent deployment, marking a significant advancement in the practical use of LLMs.
@@
While fine-tuning a small subset of parameters offers a streamlined approach for domain adaptation, it’s well-recognized that model performance is closely tied to the number of parameters involved [kaplan2020scaling]. This intrinsic characteristic of methods like LoRA often results in them falling short of the FFT baseline, which updates all parameters, thereby creating a trade-off between efficiency and model quality. This issue of compromised quality in a low-parameter setting becomes even more pronounced in target domains characterized by complex sub-domains and diverse tasks. This situation presents a compelling research question:
@@
In our research, we carry out a series of exploratory experiments, applying LoRA to the LLAMA2-7B [llama2] model to adapt it to a new domain encompassing multiple downstream tasks. As shown in Figure 1(a), LoRA adds trainable pairs of rank decomposition matrices A and B in addition to existing weight matrices. Our in-depth analysis of LoRA’s mechanics yields several insightful observations and leads to the formulation of key hypotheses. First, rather than employing a single LoRA for the entire domain, it proves more effective to deploy multiple, smaller LoRA heads, each dedicated to a specific downstream task (see Figure 1(b)). This suggests that domain or task interference might harmfully impact the training process. We further hypothesize that this interference originates from “intrinsic components”—sub-domains or distinct tasks—potentially unknown even to domain experts. Additionally, upon visualizing the parameters of LoRA, we discern a pattern: some parameters predominantly learn the commonalities across all data, while others focus on the unique aspects of each intrinsic component. From these observations, we posit that an optimal LoRA architecture should embody an explicit, asymmetric structure.
Building upon the observations, we propose an improved end-to-end LoRA framework, which we refer to as HydraLoRA. From the architecture perspective, unlike LoRA’s symmetric structure, HydraLoRA has an asymmetric structure that has a shared A matrix and multiple B matrices (see Figure 1(c)). The shared A matrix is used by all samples for parameter efficiency. During the fine-tuning phase, HydraLoRA is designed to autonomously identify “intrinsic components” and segregate training samples into distinct B matrices. During the inference phase, HydraLoRA leverages multiple B matrices using a mixture-of-experts (MoE; [JacobsJNH91, ShazeerMMDLHD17]) manner. Unlike prior work, HydraLoRA completely eliminates the need for human expertise and assumptions, showing better performance than using domain knowledge to guide the fine-tuning process.
@@ 2.1 LoRA Basics
LoRA [HuSWALWWC22] achieves comparable performances to fine-tuning on many benchmarks by freezing the pre-trained model weights and inserting trainable rank decomposition matrices into each layer of the pre-trained model. In particular, for each layer, LoRA uses two sequential low-rank matrices A and B to fit the residual weights for adaptation. The forward computation is written as follows:
$y\prime=y+\Delta y=Wx+BAx$ (1)
@@ 2.2 LoRA’s Practical Dilemma
Parameter count has a clear impact on the performance of neural models [kaplan2020scaling, gpt4]. Yet, Parameter-Efficient Fine-tuning (PEFT) methods, such as Adapter [HoulsbyGJMLGAG19] and prefix-tuning [LiL20], focus on fine-tuning a limited set of parameters. These approaches present a practical dilemma: while restricting the number of tuned parameters is essential for training efficiency, it hinders the model’s ability to learn from diverse datasets. This trade-off becomes particularly evident when considering corpus heterogeneity [S-LoRA-FL]. Figure 2 reveals a notable performance disparity between PEFT techniques and full fine-tuning (FFT), with the gap widening in scenarios involving a more diverse or heterogeneous training corpus.
Figure 2: Performance impact of corpus heterogeneity on full fine-tuning vs. parameter-efficient fine-tuning. Heterogeneity signifies the diversity within the dataset, often leading to interference due to its varied content and style [S-LoRA-FL]. Parameter-efficient approaches are particularly sensitive, suffering greater performance losses in heterogeneous cases.
2.3 Observations
In this work, we aim for a PEFT approach that strikes a better balance between maximizing the learning capability for heterogeneous data and minimizing the number of parameters involved. A key goal is to ensure that our enhanced technique exhibits robust generalization across unseen tasks, independent of any prior task-specific knowledge.
To achieve our objectives, we focus on LoRA and conduct a series of experiments as Table 1 to gain a deeper understanding of its mechanisms. Our methodology involves leveraging data from diverse tasks within a domain, and training distinct LoRA heads for each domain, leading to our first observation.
Table 1: Performance on instruction tuning with Dolly-15K [DollyV2] and evaluate with MMLU [mmlu] with different rank. For LoRA (Split) decomposes high-rank LoRA modules into smaller, equivalent low-rank components ($r\times n$). $n$ is the number of LoRAs, $r$ denotes the rank of each LoRA.
Schemes $r\times n$ MMLU $\uparrow$ % Param
@@ 2.3 Observations
This suggests that interference among tasks might harmfully impact the training process. Furthermore, we posit that this interference is NOT exclusive to this explicit multi-task training. This interference could happen in any training setting since all datasets inherently consist of multiple implicit intrinsic components, such as sub-domains or tasks within a domain that is even unknown to domain experts.
To better understand how multiple LoRA heads mitigate the interference among intrinsic components, in Figure 3, we employ the t-SNE technique [t-SNE] to visualize the parameters of matrix A and B across all heads. This analysis yields another critical observation:
(a) Compare fine-tuned LoRA modules of GSM8K [gsm8k] with its subsets using T-SNE. We employ the Independent and Identically Distributed (IID) segmentation scheme to divide GSM8K into three subsets and fine-tune them using different LoRAs.
(b) Compare fine-tuned LoRA modules of Dolly-15K [DollyV2] with its subtasks using T-SNE. We choose three subtasks of Dolly-15K including “summarization (Sum)”, “closed QA (QA)” and “information extraction (IE)”.
Figure 3: Breakdown analysis of LoRA modules. Consider LLaMA2-7b (random seed=42), which contains 32 decoder layers, corresponding to 32 adaptive modules. Each module consists of {0: q_proj of A, 1: q_proj of B, 2: v_proj of A, 3: v_proj of B} submodules. This makes a total of $32\times 4$ submodules. (a,b) left displays all submodules. (a,b) center shows all even submodules, i.e. the A matrix. (a,b) right represents all odd submodules, i.e. the B matrix. It can be seen that the differences in the fine-tuned LoRA modules for different tasks arise mainly from the B matrix.
@@
Building upon our observations, we propose an optimized LoRA architecture designed to enhance cost-effectiveness. In this architecture, we share the parameters of A matrix across various sub-domains or tasks to improve parameter efficiency, while deploying multiple B matrices, each tailored to handle different intrinsic components. This design allows for a more effective adaptation to the specific characteristics of each component. While these intrinsic components can be manually identified using prior knowledge of the training data, we also introduce end-to-end methods using K-means [k-means-1, k-means-2] and Mixture-of-Experts (MoEs) [MoE], which will be detailed in the methodology section. This automatic approach facilitates flexibility and applicability, particularly in scenarios where prior knowledge is limited or unavailable.
@@ 3 HydraLoRA
In this section, we introduce the proposed HydraLoRA, an asymmetric LoRA architecture for efficient fine-tuning, as illustrated in Figure 1. Second, we show the workflow of HydraLoRA as Figure 4.
@@ 3.1 Asymmetric LoRA architecture
The LoRA method updates two low-rank matrices A and B, and uses AB as the change of a pretrained and frozen weight W of a linear layer as shown in Eq. 1. The integral parameters are fine-tuned for the whole corpus in the original LoRA, which causes difficulty in learning the various knowledge aspects. Drawing from a detailed breakdown analysis of LoRA, a potential solution is to segment the entire LoRA into “Hydra” structured LoRA variants, that is, characterized by a central shared matrix A and several distinct matrices B, fostering a blend of shared knowledge and specialized functionalities. As Figure 1, HydraLoRA is to fine-tune LoRAs to achieve robust performance without redundancy, thereby benefiting the entire heterogeneous corpus. For the multi-head LoRA approach, we employ a set of matrices, denoted as $\mathbb{B}=\left\{\mathcal{B}_{1},\ldots,\mathcal{B}_{N}\right\}$. The asymmetric LoRA architecture can be formulated as:
$\mathbf{W}$ $=\mathbf{W}_{0}+\frac{\alpha}{r}\Delta\mathbf{W}$
$=\mathbf{W}_{0}+\frac{\alpha}{r}\sum_{j=1}^{N}\omega_{j}\cdot\mathbf{B}_{j}\mathbf{A}$
The matrics $\mathbf{B}_{i}\in\mathbb{R}^{d\times r}$ and shared $A\in\mathbb{R}^{r\times k}$. The hyper-parameter $N$ denotes the number of B modules. The term $\omega_{j}$ modulates these contribution weights for head $B_{j}$.
Figure 4: Architecture and workflow of HydraLoRA. During the fine-tuning stage, HydraLoRA first adaptively identifies and initializes $K$ of intrinsic components using $K$-means without specific domain knowledge that obtains optimal $K$ values via the SSE elbow. It then employs a trainable MoE router that treats each intrinsic component as an expert to automatically segregate training samples into intrinsic components for fine-tuning. During the merge inference stage, HydraLoRA merges multiple components flexibly and dynamically through a trainable router.
@@ 3.2 Workflow of HydraLoRA
Figure 4 illustrates the workflow of HydraLoRA. Initially, Section 3.2.1 delves into the adaptive identification and initialization of LoRA modules within a heterogeneous corpus, aligning them with task relevance through the application of $K$-means. Subsequently, we propose a Mixture-of-Experts (MoE) framework that handles LoRAs as expert adapters to ensure computational efficiency throughout the training (Section 3.2.2) and inference (Section 3.2.3) stages by freezing the rest of the LLM parameters. During inference, it flexibly and dynamically merges multiple LoRA modules through the trainable router.
3.2.1 Initialization via K-means
In the case of considering heterogeneous corpora, it is crucial to select the appropriate number $N$ of matrix B, to ensure consistent performance and minimize unnecessary computational overhead. This choice is usually closely related to the training corpus. In this work, we propose initializing HydraLoRA modules via $K$-means [k-means-1] algorithm for adaptive initialization. Specifically, $K$-means is utilized to process the heterogeneous corpus to identify the best-fit taxonomy of the corpus, i.e., the optimal $N$. First, we extract key features from the corpus by applying the Term Frequency-Inverse Document Frequency (TF-IDF; [TF-IDF]) algorithm and transform the textual information into numerical feature vectors. We integrate the elbow method [elbow] to determine the optimal value of $K$. Initially, $K$ cluster centers are randomly selected for preliminary clustering as Eq. 2, followed by updating the cluster centers to accurately reflect the data within each cluster as Eq. 3. where $C_{j}$ is the cluster center to which data point $X_{i}$ is assigned and $d(\cdot,\cdot)$ is the Euclidean distance function. $S_{j}$ is the set of data points in the $j$-th cluster.
$C_{j}$ $=\underset{C_{j}}{\mathrm{argmin}}\ d(X_{i},C_{j})$ (2)
$C_{j}$ $=\frac{1}{|S_{j}|}\sum_{X_{i}\in S_{j}}X_{i}$ (3)
3.2.2 Training
Motivated by Mixture-of-Experts (MoE; [JacobsJNH91, ShazeerMMDLHD17]), where experts are selectively activated by a gating mechanism (Router) in response to different inputs. In HydraLoRA, we substitute each expert with a lightweight LoRA adapter. During fine-tuning, while weights of LLMs remain frozen, the experts and router layers are trained from scratch. In order to achieve a unified approach to the distinct forward processes of multiple LoRAs, we define a set of experts, denoted as $(E_{1},\ldots,E_{N})$, to learn the updated matrix $\Delta W$. As HydraLoRA fine-tunes the experts using the heterogeneous corpus, the shared matrix A inherently captures collaborative knowledge to augment intra-gains, and different matrices B foster knowledge modularity to mitigate fine-tuning inter-offsets. Based on this structure, the forward process of HydraLoRA is expressed as:
$\mathbf{y}_{j}$ $=\mathbf{W}_{0}\mathbf{x}_{j}+\Delta\mathbf{W}_{j}\mathbf{x}_{j}$
$=\mathbf{W}_{0}\mathbf{x}_{j}+\sum_{i=1}^{N}\omega_{ji}\cdot E_{i}\left(\mathbf{x}_{j}\right)\quad(MoE)$ (4)
where $\mathbf{y}_{i}$ and $\mathbf{x}_{j}$ represent the input and output of intermediate LLM layers for sample $j$. $N$ denotes the number of experts, i.e., B modules, which is decided by the above K-means.To regulate these contributions, we introduce a gate function (router network) commonly consisting of a dense layer with trainable weights (transformation matrix) $\mathbf{W}_{g}\in\mathbb{R}^{r\times N}$ followed by a softmax function which takes an intermediate token representation $x$ as input and combines the output of each expert based on the gating scores $(\omega_{1},\dots,\omega_{N})$:
$\omega_{j}=\text{softmax}(\mathbf{W}_{g}^{T}x)\quad(Router)$ (5)

### v2

@@ Abstract
Adapting Large Language Models (LLMs) to new tasks through fine-tuning has been made more efficient by the introduction of Parameter-Efficient Fine-Tuning (PEFT) techniques, such as LoRA. However, these methods often underperform compared to full fine-tuning, particularly in scenarios involving complex datasets. This issue becomes even more pronounced in complex domains, highlighting the need for improved PEFT approaches that can achieve better performance. Through a series of experiments, we have uncovered two critical insights that shed light on the training and parameter inefficiency of LoRA. Building on these insights, we have developed HydraLoRA, a LoRA framework with an asymmetric structure that eliminates the need for domain expertise. Our experiments demonstrate that HydraLoRA outperforms other PEFT approaches, even those that rely on domain knowledge during the training and inference phases.
@@ 1 Introduction
Large Language Models (LLMs; [10, 3, 35, 45, 46, 31, 32]) are notably powerful, yet their training involves substantial expense. Adapting a single LLM for multiple downstream applications via fine-tuning has emerged as a prevalent method to cater to specific domain needs, balancing performance with practicality. This approach, however, faces a significant challenge due to the extensive memory and computational resources required for full fine-tuning (FFT), i.e., fine-tuning all billions of parameters. A solution to this has been the development of more selective adaptation techniques, involving modifying only a portion of the parameters or integrating external modules designed for new tasks. Key methodologies in this sphere include LoRA [17], Adaptors [36, 16, 30], and many other variants [24, 23, 9, 13, 51], all part of what can be generally termed as Parameter-Efficient Fine-tuning (PEFT). PEFT strategies are characterized by freezing the backbone model parameters while only a minimal number of task-specific parameters are introduced and fine-tuned. This method substantially boosts efficiency in the phases of fine-tuning and subsequent deployment, marking a significant advancement in the practical use of LLMs.
@@
While fine-tuning a small subset of parameters offers a streamlined approach for domain adaptation, it’s well-recognized that model performance is closely tied to the number of parameters involved [21]. This intrinsic characteristic of methods like LoRA often results in them falling short of the FFT baseline, which updates all parameters, thereby creating a trade-off between efficiency and model quality. This issue of compromised quality in a low-parameter setting becomes even more pronounced in target domains characterized by complex sub-domains and diverse tasks. This situation presents a compelling research question:
@@
In our research, we carry out a series of exploratory experiments, applying LoRA to the LLaMA2 [46] model to adapt it to a new domain encompassing multiple downstream tasks. As shown in Figure 1(a), LoRA adds trainable pairs of rank decomposition matrices A and B in addition to existing weight matrices. Our in-depth analysis of LoRA’s mechanics yields several insightful observations and leads to the formulation of key hypotheses. First, rather than employing a single LoRA for the entire domain, it proves more effective to deploy multiple, smaller LoRA heads, each dedicated to a specific downstream task (see Figure 1(b)). This suggests that domain or task interference might harmfully impact the training process. We further hypothesize that this interference originates from “intrinsic components”—sub-domains or distinct tasks—potentially unknown even to domain experts. Additionally, upon visualizing the parameters of LoRA, we discern a pattern: some parameters predominantly learn the commonalities across all data, while others focus on the unique aspects of each intrinsic component. From these observations, we posit that an optimal LoRA architecture should embody an explicit, asymmetric structure.
Building upon the observations, we propose an improved end-to-end LoRA framework, which we refer to as HydraLoRA. From the architecture perspective, unlike LoRA’s symmetric structure, HydraLoRA has an asymmetric structure that has a shared A matrix and multiple B matrices (see Figure 1(c)). The shared A matrix is used by all samples for parameter efficiency. During the fine-tuning phase, HydraLoRA is designed to autonomously identify “intrinsic components” and segregate training samples into distinct B matrices. During the inference phase, HydraLoRA leverages multiple B matrices using Mixture-of-Experts (MoE; [19, 39]) manner. Unlike prior work, HydraLoRA completely eliminates the need for human expertise and assumptions, showing better performance than using domain knowledge to guide the fine-tuning process.
@@ 2.1 LoRA Basics
LoRA [17] achieves comparable performances to fine-tuning on many benchmarks by freezing the pre-trained model weights $W_{0}$ and inserting trainable rank decomposition matrices into each layer of the pre-trained model. In particular, for each layer, LoRA uses two sequential low-rank matrices $A$ and $B$ to fit the residual weights for adaptation. The forward computation is written as follows:
$y\prime=y+\Delta y=W_{0}x+BAx$ (1)
@@ 2.2 LoRA’s Practical Dilemma
Parameter count has a clear impact on the performance of neural models [21, 32]. Yet, Parameter-Efficient Fine-tuning (PEFT) methods, such as Adapter [16] and prefix-tuning [24], focus on fine-tuning a limited set of parameters. These approaches present a practical dilemma: while restricting the number of tuned parameters is essential for training efficiency, it hinders the model’s ability to learn from diverse datasets. This trade-off becomes particularly evident when considering corpus heterogeneity [2]. Figure 2 reveals a notable performance disparity between PEFT techniques and full fine-tuning (FFT), with the gap widening in scenarios involving a more diverse or heterogeneous training corpus.
Figure 2: Performance impact of corpus heterogeneity on full fine-tuning vs. parameter-efficient fine-tuning. Heterogeneity signifies the diversity within the dataset, often leading to interference due to its varied content and style [2]. Parameter-efficient approaches are particularly sensitive, suffering greater performance losses in heterogeneous cases.
Table 1: Performance on instruction tuning with Dolly-15K [8] and evaluated with MMLU [15] with different ranks. For LoRA (Split) decomposes high-rank LoRA modules into smaller, equivalent low-rank components ($r\times n$). $n$ is the number of LoRAs, $r$ denotes the rank of each LoRA.
Schemes $r\times n$ MMLU $\uparrow$ % Parameter
@@
2.3 Observations
In this work, we aim for a PEFT approach that strikes a better balance between maximizing the learning capability for heterogeneous data and minimizing the number of parameters involved. A key goal is to ensure that our enhanced technique exhibits robust generalization across unseen tasks, independent of any prior task-specific knowledge. To achieve our objectives, we focus on LoRA and conduct a series of experiments as Table 2 to gain a deeper understanding of its mechanisms. Our methodology involves leveraging data from diverse tasks within a domain, and training distinct LoRA heads for each domain, leading to our first observation:
@@ 2.3 Observations
This suggests that interference among tasks might harmfully impact the training process. Furthermore, we posit that this interference is NOT exclusive to this explicit multi-task training. This interference could happen in any training setting since all datasets inherently consist of multiple implicit intrinsic components, such as sub-domains or tasks within a domain that is even unknown to domain experts. To better understand how multiple LoRA heads mitigate the interference among intrinsic components, in Figure 3, we employ the t-SNE technique [47] to visualize the parameters of matrix A and B across all heads. This analysis yields another critical observation:
Refer to captionRefer to captionRefer to caption
Figure 3: Breakdown analysis of LoRA modules. Compare fine-tuned LoRA modules of Dolly-15K [8] with three subtasks of Dolly-15K including “summarization (Sum)”, “closed QA (QA)” and “information extraction (IE)” using t-SNE. Consider LLaMA2-7B (random seed=42), which contains 32 decoder layers, corresponding to 32 adaptive modules. Each module consists of {0: q_proj of A, 1: q_proj of B, 2: v_proj of A, 3: v_proj of B} submodules. This makes a total of $32\times 4$ submodules. Left displays all submodules. Center shows all even submodules, i.e. the A matrix. Right represents all odd submodules, i.e. the B matrix. It can be seen that the differences in the fine-tuned LoRA modules for different tasks arise mainly from the B matrix.
@@
Building upon our observations, we propose an optimized LoRA architecture designed to enhance cost-effectiveness. In this architecture, we share the parameters of A matrix across various sub-domains or tasks to improve parameter efficiency, while deploying multiple B matrices, each tailored to handle different intrinsic components. This design allows for a more effective adaptation to the specific characteristics of each component. While these intrinsic components can be manually identified using prior knowledge of the training data, we also introduce end-to-end methods using Mixture-of-Experts (MoEs) [20], which will be detailed in the methodology section. This automatic approach facilitates flexibility and applicability, particularly in scenarios where prior knowledge is limited or unavailable.
@@ 3 HydraLoRA
In this section, we introduce the proposed HydraLoRA, an asymmetric LoRA architecture for efficient fine-tuning, as illustrated in Figure 1. After that, we show the workflow of HydraLoRA as Figure 4.
@@ 3.1 Asymmetric LoRA architecture
The LoRA method updates two low-rank matrices $A$ and $B$, and uses $AB$ as the change of a pretrained and frozen weight $W_{0}$ of a linear layer as shown in Eq. 1. The integral parameters are fine-tuned for the whole corpus in the original LoRA, which causes difficulty in learning the various knowledge aspects. Drawing from a detailed breakdown analysis of LoRA, a potential solution is to segment the entire LoRA into “Hydra” structured LoRA variants, that is, characterized by a central shared matrix $A$ and several distinct matrices $B$, fostering a blend of shared knowledge and specialized functionalities. As Figure 1, HydraLoRA is to fine-tune LoRAs to achieve robust performance without redundancy, thereby benefiting the entire heterogeneous corpus. The asymmetric LoRA architecture can be formulated as:
$\mathit{W}$ $=\mathit{W_{0}}+\Delta\mathit{W}$ (2)
$=\mathit{W_{0}}+\sum_{i=1}^{N}\omega_{i}\cdot B_{i}A$
The matrics $B_{i}\in\mathbb{R}^{d\times r}$ and shared $A\in\mathbb{R}^{r\times k}$. The hyper-parameter $N$ denotes the number of $B$ matrices. The term $\omega_{i}$ modulates these contribution weights for head $B_{i}$.
Figure 4: Architecture and workflow of HydraLoRA. During the fine-tuning stage, HydraLoRA first adaptively identifies and initializes $k$ of intrinsic components without specific domain knowledge. It then employs a trainable MoE router that treats each intrinsic component as an expert to automatically segregate training samples into intrinsic components for fine-tuning. During the inference stage, HydraLoRA merges multiple $B$ matrices flexibly and dynamically through a trained router.
@@ 3.2 Workflow of HydraLoRA
Figure 4 illustrates the workflow of HydraLoRA. Initially, HydraLoRA delves into the adaptive identification and initialization of LoRA modules within a heterogeneous corpus, aligning them with task relevance through the application of $k$-means or developer-specified size. Subsequently, we propose a Mixture-of-Experts (MoE) framework that handles $B$ matrices as expert adapters to ensure computational efficiency throughout the fine-tuning (Section 3.2.1) and inference (Section 3.2.2) stages by freezing the rest of the LLM parameters. During inference, it flexibly and dynamically merges multiple $B$ matrices through the MoE router.
3.2.1 Fine-tuning
Motivated by Mixture-of-Experts (MoE; [19, 39]), where experts are selectively activated by a gating mechanism (Router) in response to different inputs. In HydraLoRA, we substitute each expert with a lightweight LoRA adapter. During fine-tuning, while weights of LLMs remain frozen, the experts and router layers are trained from scratch. In order to achieve a unified approach to the distinct forward processes of multiple $B$ matrices, we define a set of experts, denoted as $(E_{1},\ldots,E_{N})$, to learn the updated matrix $\Delta W$. As HydraLoRA fine-tunes the experts using the heterogeneous corpus, the shared matrix $A$ inherently captures collaborative knowledge to augment intra-gains, and different matrices $B$ foster knowledge modularity to mitigate fine-tuning inter-offsets. Based on this structure, the forward process of HydraLoRA is expressed as:
$y=\mathit{W}_{0}x+\sum_{i=1}^{N}\omega_{i}E_{i}Ax\quad(MoE)$ (3)
where $N$ denotes the number of experts, i.e., $B$ matrices.To regulate these contributions, we introduce a gate function (router network) commonly consisting of a dense layer with trainable weights (transformation matrix) $\mathit{W}_{g}\in\mathbb{R}^{r\times N}$ followed by a softmax function which takes an intermediate token representation $x$ as input and combines the output of each expert based on the gating scores $(\omega_{1},\dots,\omega_{N})$:
$\omega_{i}=\text{softmax}(\mathit{W}_{g}^{T}x)\quad(Router)$ (4)